[Hallc_running] [New Logentry] RC daily update (8/26)

billlee at jlab.org billlee at jlab.org
Wed Aug 26 09:35:01 EDT 2020

Logentry Text:
Dear Collaborators,

I am happy to report that we entered the production phase with He3 target started around midnight. 

In the past 24 hours: the accelerator went down for the planned 2 hour maintenance yesterday at 8:30am to repair 1L13-7 cavity, and came back at round 12:00. Junhao took an opportunistic access into the hall and  brought back the EPR system, and he was able to perform the EPR calibration before the moller measurement. Moller started at 12:00 and completed 4:00pm, it was very smooth. We completed one more setting of the reference cell data: Ref-H2 at Kin C-Z. The NMR measurement at 23:50 was 45.2% (with the real calibration constants).   

Overnight, we ran into a few issues with the beam stability (Hall C ion chamber tripping due to charge build-up even without beam) and current reduction (we requested 30uA but only got ~24uA), the accelerator crews are looking into the possible solutions for help with these issue. Throughout today, we aim to run production, while detector experts will continue optimizing the target polarization. 
Please take a look at the most recent progress at: https://hallcweb.jlab.org/wiki/index.php/RC_Update_and_Run_Plan_for_Aug_26,_2020

We will have 4:15pm meeting today. 
Meeting link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3834786
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200826/d300a571/attachment.html>

More information about the Hallc_running mailing list